### v1

@@ 1 Introduction
Recently, large pre-trained vision-language (V-L) models like CLIP [51] have achieved huge success on various fields. The excellent generalization ability is a crucial cornerstone of such success. Decent results could be reached without adaptation on downstream tasks [51, 22, 59, 52, 46].
When people have access to downstream data, it is better to tune the pre-trained V-L models on the target dataset for higher accuracy. However, such full fine-tuning could easily cause the model to overfit the small downstream dataset and face catastrophic forgetting problem, leading to severe performance drop.
In this paper, we will start from rethinking the relationship between noise and generalization. We propose to think noise as a kind of regularization techniques, which may be helpful for alleviating overfitting and catastrophic forgetting problem. As in Sec. 3.2 and Fig. 2, we then find that directly adding some random Gaussian noise to the tunable prompts of vision-language models would result in a performance gain in several cases, which partly validates our conjecture about using noise as a form of regularization. In particular, excessive noise diminishes the model’s adaptation capability, while insufficient noise fails to provide effective regularization. Only noise of moderate intensity is beneficial for the model’s generalization.
@@
Since our design is general, we could fruitfully integrate our quantization strategy into many existing methods and reach a higher accuracy with much smaller model size due to the lightweight nature of quantization. For example, we integrate our method onto an existing popular method MaPLe, and our QMaPLe earns 0.57% accuracy improvement with only 0.25×0.25$\times$0.25 × size comapred with MaPLe.
@@ 2.2 Parameter-Efficient Fine-Tuning
Parameter-efficient fine-tuning (PEFT) [20, 19, 40, 34, 23, 68, 67, 37, 16, 25, 18, 57] aims to transfer a pre-trained model to numerous downstream tasks using a minimum number of parameters. Originating from natural language processing tasks, classic methods like adapter [19], prompt tuning [34, 31, 39, 54, 38, 24] and LoRA [20] follows a similar principle to add extra modules with a small number of parameters into the backbone model, freeze the original parameters and only tune and store the newly added parameters. Inspired by the success of PEFT in language field, researchers have extended such kind of approaches to adapt visual models in a similar fashion [23, 65, 16, 37, 6]. In the field of vision-language modeling, several explorations have been made as well. Bahng et al.[1] exclusively apply prompt tuning to the image encoder. CoOp[68] replaces the fixed template in CLIP [51] with tunable text prompts. CoCoOp [67] leverages image features to guide the optimization of tunable text prompts in CoOp. Other works [28, 30] optimize both image and text prompts simultaneously and establish additional connections between different modalities. To mitigate overfitting and catastrophic forgetting, various works [29, 58, 5, 66] integrate regularization modules or losses into prompt tuning.
@@ 5 Experiments
CLIP 0KB 69.34 74.22 71.70
CoOp 4.1KB 82.69 63.22 71.66
CoCoOp 70.8KB 80.47 71.69 75.83
Adapter 1051KB 82.62 70.97 76.35
LoRA 258KB 84.30 67.33 74.86
ProGrad 16.4KB 82.79 68.55 75.00

### v2

@@ 1 Introduction
Recently, deep learning models and related technologies have seen rapid development [23, 51, 50, 67, 66, 13, 24, 25, 12, 11, 9]. Vision-language model (VLM) like CLIP [59] is one of the hottest research topics and leads to huge success. The excellent generalization ability is a crucial cornerstone of such achievements [59, 27, 69, 60, 54].
When people have access to downstream data, it is better to tune the pre-trained VLM on the target dataset for higher accuracy. However, such full fine-tuning could easily cause the model to overfit the small downstream dataset and face catastrophic forgetting problem, leading to severe performance drop.
To solve the problem, in this paper we will start from rethinking the relationship between noise and generalization. We propose to think noise as a kind of regularization techniques, which may be helpful for alleviating overfitting and catastrophic forgetting problem. As in Sec. 3.2 and Fig. 2, we then find that directly adding some random Gaussian noise to the tunable prompts of vision-language models would result in a performance gain in several cases, which partly validates our conjecture about using noise as a form of regularization. In particular, excessive noise diminishes the model’s adaptation capability, while insufficient noise fails to provide effective regularization. Only noise of moderate intensity is beneficial for the model’s generalization.
@@
Since our design is general, we could fruitfully integrate our quantization strategy into many existing methods and reach a higher accuracy with much smaller model size due to the lightweight nature of quantization. For example, we integrate our method onto an existing popular method MaPLe, and our QMaPLe earns 0.57% accuracy improvement with only 0.25×0.25$\times$0.25 × size.
@@ 2.2 Parameter-Efficient Fine-Tuning
Recently, a series of works [25, 24, 45, 39, 28, 78, 77, 42, 20, 30, 22, 65, 49] have been proposed to help transfer the learnt knowledge, where one of the most popular field is parameter-efficient fine-tuning (PEFT). PEFT aims to transfer a pre-trained model to downstream tasks by a minimum number of parameters. Originating from natural language processing tasks, classic methods like adapter [24], prompt tuning [39, 36, 44, 62, 43, 29] and LoRA [25] follows a similar principle to add extra modules with a small number of parameters into the backbone model, freeze the original parameters and only tune and store the newly added parameters. Inspired by the success of PEFT in language field, researchers have extended such kind of approaches to adapt visual models in a similar fashion [28, 75, 20, 42, 6]. In the field of vision-language modeling, several explorations have been made as well. Bahng et al.[1] exclusively apply prompt tuning to the image encoder. CoOp[78] replaces the fixed template in CLIP [59] with tunable text prompts. CoCoOp [77] leverages image features to guide the optimization of tunable text prompts in CoOp. Other works [33, 35] optimize both image and text prompts simultaneously and establish additional connections between different modalities. To mitigate overfitting and catastrophic forgetting, various works [34, 68, 5, 76] integrate regularization modules or losses into prompt tuning.
@@ 5 Experiments
CLIP [59] 0KB 69.34 74.22 71.70
CoOp [78] 4.1KB 82.69 63.22 71.66
CoCoOp [77] 70.8KB 80.47 71.69 75.83
Adapter [17] 1051KB 82.62 70.97 76.35
LoRA [25] 258KB 84.30 67.33 74.86
ProGrad [79] 16.4KB 82.79 68.55 75.00
@@ Acknowledgements.
Appendix 0.A Datasets
Building upon prior research [78, 77], we utilize eleven datasets pertaining to image recognition to substantiate the effectiveness of the proposed methodology in addressing the base-to-new generalization task. These datasets encompass two repositories dedicated to generic object classification, namely ImageNet and Caltech101, five repositories catering to fine-grained classification, including OxfordPets, StanfordCars, Flowers102, Food101, and FGVCAircraft, one repository for scene recognition, denoted as SUN397, one repository for action recognition, known as UCF101, one repository for texture classification, termed DTD, and one repository for satellite imagery recognition, designated EuroSAT. Consistent with earlier studies [78, 77, 33, 79], for each dataset in base-to-new generalization, we evenly partition the classes into two distinct groups that do not overlap, with one group serving as the base classes and the other as the new classes. We train all models only using the base classes and conduct evaluation on both the base and new classes to verify the specialization capability and generalization capability of the models.
In the domain generalization task, we leverage ImageNet-A, ImageNet-R, ImageNetv2, and ImageNet-S to assess the robustness of the model. In this context, the model is initially trained using ImageNet, followed by direct utilization of images from the aforementioned datasets for inference.
Concerning the cross-dataset transfer task, the datasets mirror those utilized in the base-to-new generalization task. Analogous to domain generalization, the model undergoes initial training on ImageNet followed by inference on the remaining ten distinct datasets.
For the few-shot learning task, the datasets align with those employed in the base-to-new generalization task. The model is trained and assessed with varying numbers of shots, specifically 1, 2, 4, 8, and 16 shots separately.
The dataset partitioning mirrors that of earlier works [78, 77]. We present the average model performance over three iterations with distinct random seeds to ensure fair comparisons.
Table 8: Full results in base-to-new generalization. H: harmonic mean [64].
(a) An overview of the size of different methods..
Method CoOp CoCoOp Adapter LoRA ProGrad QCoOp MaPLe QMaPLe
size 4.1KB 70.8KB 1051KB 258KB 16.4KB 0.26KB 7096KB 1774KB
(b) Average
Base New H
CLIP 69.34 74.22 71.70
CoOp 82.69 63.22 71.66
CoCoOp 80.47 71.69 75.83
Adapter 82.62 70.97 76.35
LoRA 84.30 67.33 74.86
ProGrad 82.79 68.55 75.00
QCoOp 80.68 74.44 77.43
MaPLe 82.28 75.14 78.55
QMaPLe 83.02 75.57 79.12
(c) ImageNet
Base New H
CLIP 72.43 68.14 70.22
CoOp 76.47 67.88 71.92
CoCoOp 75.98 70.43 73.10
Adapter 76.53 66.67 71.26
LoRA 74.77 58.47 65.62
ProGrad 77.03 68.80 72.68
QCoOp 76.17 70.73 73.35
MaPLe 76.66 70.54 73.47
QMaPLe 76.93 70.73 73.70
(d) Caltech101
Base New H
CLIP 96.84 94.00 95.40
CoOp 98.00 89.81 93.73
CoCoOp 97.96 93.81 95.84
Adapter 98.20 93.20 95.63
LoRA 98.49 90.33 94.24
ProGrad 98.50 91.90 95.09
QCoOp 97.80 95.03 96.40
MaPLe 97.74 94.36 96.02
QMaPLe 97.97 95.00 96.46
(e) OxfordPets
Base New H
CLIP 91.17 97.26 94.12
CoOp 93.67 95.29 94.47
CoCoOp 95.20 97.69 96.43
Adapter 94.40 94.10 94.25
LoRA 94.90 92.57 93.72
ProGrad 94.40 95.10 94.75
QCoOp 95.17 97.60 96.37
MaPLe 95.43 97.76 96.58
QMaPLe 95.67 97.63 96.64
(f) StanfordCars
Base New H
CLIP 63.37 74.89 68.65
CoOp 78.12 60.40 68.13
CoCoOp 70.49 73.59 72.01
Adapter 77.13 69.23 72.97
LoRA 81.07 65.30 72.34
ProGrad 79.00 67.93 73.05
QCoOp 73.73 72.90 73.31
MaPLe 72.94 74.00 73.47
QMaPLe 75.00 73.67 74.33
(g) Flowers102
Base New H
CLIP 72.08 77.80 74.83
CoOp 97.60 59.67 74.06
CoCoOp 94.87 71.75 81.71
Adapter 97.70 70.83 82.13
LoRA 98.23 60.20 74.65
ProGrad 96.27 71.07 81.77
QCoOp 95.57 74.67 84.22
MaPLe 95.92 72.46 82.56
QMaPLe 96.43 74.33 83.95
(h) Food101
Base New H
CLIP 90.10 91.22 90.66
CoOp 88.33 82.26 85.19
CoCoOp 90.70 91.29 90.99
Adapter 90.40 90.40 90.40
LoRA 88.57 87.30 87.93
ProGrad 90.17 89.53 89.85
QCoOp 90.87 91.90 91.38
MaPLe 90.71 92.05 91.38
QMaPLe 90.63 92.10 91.36
(i) FGVCAircraft
Base New H
CLIP 27.19 36.29 31.09
CoOp 40.44 22.30 28.75
CoCoOp 33.41 23.71 27.74
Adapter 39.57 32.27 35.55
LoRA 46.27 28.83 35.53
ProGrad 42.63 26.97 33.04
QCoOp 37.50 34.03 35.68
MaPLe 37.44 35.61 36.50
QMaPLe 39.10 34.90 36.88
(j) SUN397
Base New H
CLIP 69.36 75.35 72.23
CoOp 80.60 65.89 72.51
CoCoOp 79.74 76.86 78.27
Adapter 81.67 73.93 77.61
LoRA 79.73 69.00 73.98
ProGrad 80.70 71.03 75.56
QCoOp 79.20 77.93 78.56
MaPLe 80.82 78.70 79.75
QMaPLe 81.33 78.27 79.77
(k) DTD
Base New H
CLIP 53.24 59.90 56.37
CoOp 79.44 41.18 54.24
CoCoOp 77.01 56.00 64.85
Adapter 80.47 52.23 63.35
LoRA 82.93 54.90 66.06
ProGrad 76.70 46.67 58.03
QCoOp 74.97 58.37 65.63
MaPLe 80.36 59.18 68.16
QMaPLe 80.77 57.63 67.27
(l) EuroSAT
Base New H
CLIP 56.48 64.05 60.03
CoOp 92.19 54.74 68.69
CoCoOp 87.49 60.04 71.21
Adapter 86.93 64.20 73.86
LoRA 94.90 65.67 77.62
ProGrad 91.37 56.53 69.85
QCoOp 83.53 69.80 76.05
MaPLe 94.07 73.23 82.35
QMaPLe 94.30 79.47 86.25
(m) UCF101
Base New H
CLIP 70.53 77.50 73.85
CoOp 84.69 56.05 67.46
CoCoOp 82.33 73.45 77.64
Adapter 85.80 73.63 79.25
LoRA 87.47 68.03 76.53
ProGrad 83.90 68.50 75.42
QCoOp 81.87 75.93 78.79
MaPLe 83.00 78.66 80.77
QMaPLe 85.10 77.50 81.12
Appendix 0.B Training Configuration
Following the conventional setup outlined in [77], we employ ViT-B/16 as the image encoder within CLIP. Prior to feeding into the image encoder, each training image is resized to $224\times 224$. To augment the data, standard techniques such as random cropping and flipping are applied, consistent with the methodology described in [77]. During training, a batch size of 32 is utilized, and stochastic gradient descent (SGD) is employed to optimize the learnable parameters. Similar to the approach detailed in [78], a warm-up scheme is implemented during the first epoch, which proves crucial for prompt tuning. All other baselines are configured strictly according to the specifications provided in their respective original papers.
Hyperparameter tuning is performed via a grid search methodology, guided by the parameter configurations reported in previous studies [78, 33]. For experiments involving QCoOp, the quantization bit of the prompts is set to 1 by default unless otherwise specified. In the case of QMaPLe experiments, the parameters in the projection layer, responsible for transforming text prompts into image prompts, and the parameters in the prompts are all subjected to quantization, with a quantization bit of 4. Additionally, following the structure of MaPLe, nine transformer layers are typically modified within QMaPLe experiments.
Table 9: Results of deep prompts for QCoOp.
total depth size base new H
1 0.26KB 79.49 72.65 75.92
2 0.52KB 78.19 73.67 75.86
3 0.78KB 79.82 72.27 75.86
4 1.04KB 80.77 72.31 76.31
5 1.30KB 81.14 72.67 76.67
6 1.56KB 81.65 72.68 76.90
Appendix 0.C Full Base-To-New Generalization Results
As shown in Tab. 8, QCoOp and QMaPLe significantly improve the generalization capability represented by the accuracy on the new classes. Compared with CoOp, QCoOp earns $11.22$% accuracy gain on the new classes and $2.10$% accuracy drop on the base classes. Among all the lightweight SOTA methods, QCoOp gets strongest harmonic mean accuracy on 7 out of 11 datasets including ImageNet, Caltech101, StanfordCars, Flowers102, Food101, FGVCAircraft, and SUN397. Compared with a heavy method MaPLe, QMaPLe achieves better harmonic mean accuracy on 9 out of 11 datasets, including ImageNet, Caltech101, OxfordPets, StanfordCars, Flowers102, FGVCAircraft, SUN397, EuroSAT and UCF101.
Appendix 0.D Experiments on Deep Prompt Configurations
In this paragraph, we investigate the impact of deep prompts, as described in Eq. 4 of the main text. By default, QCoOp employs prompts with a depth of 1, which means the prompts are only added to the input layer. We systematically increase the number of tunable prompts in subsequent transformer layers. Considering that the text transformer in CLIP consists of 12 layers, we add prompts to a maximum of 6 layers for verification. As illustrated in Table Tab. 9, incorporating more prompts across multiple layers tends to improve the specialized capability, albeit at the expense of a slight reduction in generalized capability. In summary, at the expense of increased dimensionality, using more prompts across multiple layers can improve the harmonic mean accuracy to some extent.
Table 10: Comparisons of using quantized prompts in textual encoder and image encoder at minimal size.
size base new H
QCoOp 0.26KB 79.49 72.65 75.92
QVPT 0.39KB 73.15 68.93 70.98
Appendix 0.E Prompt Modality Considerations for Minimizing Storage Cost
In this paragraph, we conduct a brief comparison of quantizing prompts in the visual transformer, following the approach outlined in VPT [28]. The results are presented in Tab. 10. It is evident that when operating under a stringent storage constraint, QCoOp outperforms QVPT in terms of both accuracy and size efficiency. Adding prompts to the textual transformer is better.